a collection of 5 centavo coins, 10 centavo coins and 25 centavo coins totals P4.20. The total number of coins is 36. There are twice as many 10 centavo coins as the 25 centavo coins. How many of each kind are there??? PLEASE HELP ME...I REALLY NEED TO ANSWER MY HOMEWORK...BECAUSE THESE ARE ALL PART OF OUR FINAL EXAMS..THANX...
Answer by stanbon(75887)

a collection of 5 centavo coins, 10 centavo coins and 25 centavo coins totals P4.20. The total number of coins is 36. There are twice as many 10 centavo coins as the 25 centavo coins. How many of each kind are there???
--------------------------------------------
Let the number of 25 cent coins be "x" ; value is 25x cents
Number of 10 cent coins is "2x"; value is 10(2x) = 20x cents
Number of 5 cent coins is (36-x-2x) ; value is 5(36-3x) = 180-15x cents
=====================
EQUATION:
value + value + value = 420 cents
25x + 20x + 180-15x = 420
30x = 240
x = 8 (# of 25 cent coins)
2x = 16 (# of 10 cent coins)
36-3x = 36-24 = 12 (# of 5 cent coins)
=======================
Cheers,
Stan H.